# Multi-Level Explanations for Generative Language Models

### v1

@@ 5 User Study
We conducted a user study to understand how humans perceive explanations provided by different scalarizers and attribution methods, and whether they can discern performance differences akin to the quantitative evaluations in Section 4. To ease interpretation, we developed a visualization tool that highlights input text spans based on a color-coded scale for the attribution scores. We addressed the following research questions for summarization:
@@
  3. 3.
    Concentration: What are people’s perceptions regarding the concentration of attribution scores (i.e., do the darker colors cover less of the text)?
  4. 4.
    Granularity: Which granularity level (sentence-level vs. multi-level) do people prefer?
We selected ten examples from the XSUM dataset for the user study with diversity in topics, while ensuring that they do not contain sensitive issues or obvious errors. We designed an online survey consisting of three parts. Each part showed an input text, randomly drawn from the ten examples, and its summary, generated by the DistilBart model. This user study focused on algorithms and scalarizers that are of most interest based on the automated evaluation in Section 4. Specifically, we compared two scalarizers (Log Prob, BERT) in Part 1, three attribution algorithms (C-LIME, L-SHAP, PartitionSHAP) in Part 2, and two levels of granularity (sentence-level, multi-level) in Part 3. The presentation order of the attribution algorithms was randomized to mitigate order effects. The survey consisted of seven pairwise comparisons in total followed by questions for the participants.
We recruited participants from a large technology company who self-identify as machine learning practitioners using language models and collected data from 88 of them after filtering. Here, we report a summary of key results only. See Appendix D for details including survey questions, analysis, and statistical results.
Scalarizers. Significantly more participants perceived BERT to be higher in fidelity than Log Prob. They also preferred BERT over Log Prob. Participants perceived that the attribution concentration was adequate in both BERT and Log Prob.
Attribution methods. Significantly more participants perceived C-LIME to be higher in fidelity than L-SHAP. They also preferred C-LIME over L-SHAP. Participants perceived that the attribution concentration was adequate in all three methods.
Granularity. Regarding user preference between sentence-level and multi-level granularities, no statistically significant difference was found.
@@ 6 Concluding Remarks
We proposed MExGen, a framework to provide explanations for generative language models. Our experiments show that instances of MExGen provide faithful explanations and outperform the baselines we compare against — PartitionSHAP and CaptumLIME. Our numerical results also indicate that C-LIME outperforms the other explanation methods in the MExGen framework and PartitionSHAP. The user study results align with the automated evaluation overall, and additionally reveal that people perceive the BERT scalarizer as more locally faithful than the Log Prob scalarizer, which was not seen in the automated evaluation. This result implies that in some cases, there may be no loss in having text-only access compared to full logit access.
@@
We see three main limitations in our work. First, MExGen is a framework for post hoc explanations. Although providing such explanations can be helpful for practitioners to understand the model behavior, it does not fully characterize how models make predictions and only provides local explanations. Second, our automatic evaluations use three models (DistilBart, Flan-UL2, and Flan-T5-Large) and three datasets (XSUM, CNN/DM, and SQUaD). Although our findings are consistent across the models and datasets, the results reported in Section 4 could still change in different experimental settings. Third, our user study analyzes the perception of participants of how well a method explains the predictions of a given model, and not necessarily the fidelity of the explanation itself — fidelity is measured more directly in the automated evaluation. However, we believe the fact that the participant pool was composed of people with experience in ML and LLMs improves the quality of evaluation of the explanation methods. Finally, with post hoc explanations, there is the risk that the explanation algorithm could be steered to obfuscate undesirable behavior from the model.
@@ Acknowledgements
The authors thank the following people: Inkit Padhi for suggestions on parsing sentences into phrases, using language models for perturbing words, and participation in multiple discussions; Ella Rabinovich and Samuel Ackerman for discussions on and recommendations of similarity measures as scalarizers, and Ella in particular for also providing suggestions on parsing sentences into phrases and using language models for perturbation; Pratap Kishore Varma Vemulamanda for helping to investigate the API-only case of SHAP’s summarization example [24]; Subhajit Chaudhury for suggestions of similarity measures as scalarizers; Keshav Ramji for discussion on related work and quality evaluation for the explanations; Kush Varshney and Eitan Farchi for general encouragement and support.
Appendix A Further Details on MExGen
A.1 Scalarizers
@@
The inputs $x$ in our case are perturbed versions of the original input $x^{o}$. Some of these perturbations however may be significantly different semantically and cause the probability of generating the original output $y^{o}$ to decrease dramatically. For this reason, the Log Prob scalarizer may require access to logits for improbable tokens conditioned on $x$, which can be ensured if logits are available for the entire vocabulary, but not if only the top $k$ logits are provided.
@@
[…]
@@
This scalarizer is based on the intuition that Log NLI entailment is a kind of similarity. We use the NLI model to predict the log-odds of entailment given $y^{o}$ as premise and $y$ as hypothesis, and optionally in the other direction as well. If both directions are used, we take the geometric mean of the two entailment probabilities and then convert back to log-odds.
A.2 Phrase Segmentation
Here we describe the phrase segmentation algorithm mentioned in Section 3.2. The algorithm starts with the root token of the tree and checks whether each child subtree of the root is shorter (in terms of tokens) than a maximum phrase length parameter. If it is shorter, then the subtree constitutes a phrase, and if it is not, the algorithm is recursively applied to the subtree. The root of each (sub)tree is also taken to be a phrase. Once the sentence has been recursively segmented into phrases in this manner, a second pass is performed to re-merge some phrases that have become too fragmented, thus controlling the number of phrases which is desirable for computation and interpretation. Specifically, phrases that constitute noun chunks (as identified by spaCy) are merged, and certain single-token phrases are merged with their neighbors. Further notes:
@@ A.2 Phrase Segmentation
A.3 C-LIME
In addition to the departures from LIME described in Section 3.2, we discuss two other aspects:
Sample weighting
Since we limit the number of units that are simultaneously perturbed to a small integer $K$, we also no longer use LIME’s sample weighting scheme. Instead, we give each subset cardinality $k=0,\dots,K$ the same total weight, and then distribute this weight uniformly over the subsets of that cardinality.
Regularization
Different regularizers $R(w)$ can be used in (3), e.g. $\ell_{2}$ or $\ell_{1}$. In our experiments however, we do not regularize and compute a fully dense solution. This allows all units to be ranked to facilitate evaluation of perturbation curves.
Appendix B Algorithm Parameter Settings
This appendix documents the settings of key algorithm parameters and other choices made in the experiments.
@@ Appendix B Algorithm Parameter Settings
  • Sim”: We use the all-MiniLM-L6-v2 embedding model from the SentenceTransformers package666this model can be found at https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2.
  • BERT”: We use the Log NLI model deberta-v2-xxlarge-mnli777this model can be found at https://huggingface.co/microsoft/deberta-v2-xxlarge-mnli [7] to compute BERTScore, which is the same Log NLI model that we use for the Log NLI scalarizer. Our initial reason for doing so was to see whether the two scalarizers would be very similar because of this choice (they are not as they operate on different principles). We take the F1-score output as the BERTScore.
  • BART” and “SUMM”: For “SUMM”, we follow SHAP [24] in using the same distilbart-xsum-12-6 model as both the scalarizing summarization model as well as the primary summarization model to explain. For “BART”, we also use distilbart-xsum-12-6 to determine whether the “BART” and “SUMM” scalarizers are very similar when instantiated with the same model. This is indeed the case.
@@
The phrase segmentation algorithm has one main parameter, the maximum phrase length, which we set to $10$ spaCy tokens (not counting spaces and punctuation). We exclude non-alphanumeric units (generally punctuation) from attribution. For QA, we allow only the context to be attributed to.
@@
The decision here is how many top sentences are refined into phrases in the summarization experiments and into words in the QA experiments. For summarization, this is determined by first normalizing the sentence-level attribution scores to the range $[-1,1]$ and then choosing sentences with absolute attribution scores above a threshold. The threshold is $1/3$ for C-LIME and LOO with the distilbart-xsum-12-6 model, $0.3$ for L-SHAP with distilbart-xsum-12-6, $0.5$ for C-LIME and LOO with the Flan-UL2 model, and $0.4$ for L-SHAP with Flan-UL2. If more than three sentences have scores above the threshold, we take only the top three. For QA, we simply take the top $1$ sentence as the context paragraphs in SQuAD tend to have only a handful of sentences.
@@
For L-SHAP, the two main parameters are the local neighborhood radius $M$, which we take to be $M=2$, and the maximum number of units $K$ perturbed at one time, also set to $2$. Note that the latter does not include the unit of interest, so altogether the maximum is $3$.
C-LIME
For C-LIME, the main parameters controlling the perturbations are the constant of proportionality between the number of perturbations $n$ and number of units $d$, and the maximum number of units $K$ perturbed at one time. The former is set to $10$ and the latter to $3$, except for sentence-level attributions on SQuAD where it is $2$.

### v2

@@ 5 User Study
We conducted a user study to understand how humans perceive explanations provided by different scalarizers and attribution methods, and whether they can discern performance differences akin to the quantitative evaluations in Section 4. To ease interpretation, we developed a visualization tool that highlights input text spans based on a color-coded scale for the attribution scores. In this section, we focus on the following research questions:
@@
We discuss additional research questions on concentration and granularity of attribution scores in Appendix E.
We selected ten examples from the XSUM dataset for the user study with diversity in topics, while ensuring that they do not contain sensitive issues or obvious errors. We designed an online survey consisting of three parts. Each part showed an input text, randomly drawn from the ten examples, and its summary, generated by the DistilBART model. This user study focused on algorithms and scalarizers that are of most interest based on the automated evaluation in Section 4. Specifically, we compared two scalarizers (Log Prob, BERT) and three attribution algorithms (C-LIME, L-SHAP, PartitionSHAP). The presentation order of the attribution algorithms was randomized to mitigate order effects. The survey consisted of seven pairwise comparisons in total followed by questions for the participants.
We recruited participants from a large technology company who self-identify as machine learning practitioners using language models and collected data from 88 of them after filtering. Here, we report a summary of key results only. See Appendix E for details including survey questions, analysis, and statistical results.
Scalarizers. Significantly more participants perceived BERT to be higher in fidelity than Log Prob (57% to 35%). They also preferred BERT over Log Prob (64% to 31%). This result is notable because BERT uses only text output from the LM while Log Prob depends on output logits.
Attribution methods. Significantly more participants perceived C-LIME to be higher in fidelity than L-SHAP ($p$-value $=0.011$). They also preferred C-LIME over L-SHAP ($p=0.007$). This result is notable because C-LIME and L-SHAP performed very similarly in the automated evaluation in Section 4.3.
@@ 6 Concluding Remarks
We proposed MExGen, a framework to provide explanations for LLMs used in context-grounded tasks like summarization and question answering. MExGen uses a multi-level strategy to efficiently explain model predictions in the case of long inputs. MExGen can produce explanations even when only text outputs are available (API access), thanks to scalarizers that map text to numerical values. Our experiments offer guidance on the choice of scalarizer and show that instances of MExGen provide more faithful explanations, outperforming the baselines PartitionSHAP and CaptumLIME as well as self-explanations by powerful LLMs. The user study results align with the automated evaluation, and reveal that people perceive the BERT scalarizer as more locally faithful than the Log Prob scalarizer. This result implies that in some cases, there may be no loss in having text-only access compared to full logit access.
@@
We see the following limitations and risks:
(1) MExGen is a framework for post hoc explanations. Although such explanations can help practitioners understand model behavior, they do not fully characterize how models generate output and only provide local explanations.
(2) Although the findings of our automated evaluation are consistent across the tested models and datasets, the results reported in Section 4 could still change in other experimental settings.
(3) Our user study analyzes the perception of participants of how well a method explains the predictions of a model, and not necessarily the fidelity of the explanation itself — fidelity is measured more directly in the automated evaluation. However, we believe the fact that the participant pool was composed of people with experience in ML and LLMs improves the quality of their evaluation.
(4) Post hoc explanations in general come with the risk of being steered to obfuscate undesirable behavior from the model. One potential mitigation is to couple them with additional types of explanation, where possible. Another mitigation is to have a different party compute explanations, rather than the model developer. The black-box, perturbation-based explanations considered in this work lend themselves to such a two-party scenario, where the party computing explanations only needs to query the model to do so. This party would select which perturbed inputs to query the model on, avoiding one path to obfuscation where a model developer uses specially chosen perturbations to conceal undesirable behavior. The explaining party should also be given a large budget of model queries to better probe model behavior.
@@ Acknowledgements
The authors thank the following people: Inkit Padhi for suggestions on parsing sentences into phrases, using language models for perturbing words, and participation in multiple discussions; Ella Rabinovich and Samuel Ackerman for discussions on and recommendations of similarity measures as scalarizers, and Ella in particular for also providing suggestions on parsing sentences into phrases and using language models for perturbation; Pratap Kishore Varma Vemulamanda for helping to investigate the API-only case of SHAP’s summarization example SHAP (2024b); Subhajit Chaudhury for suggestions of similarity measures as scalarizers; Keshav Ramji for discussion on related work and quality evaluation for the explanations; Kush Varshney and Eitan Farchi for general encouragement and support. The authors also thank the anonymous ARR reviewers and action editor for their high-quality reviews and positive, constructive comments, especially Reviewer GEAE for suggesting the comparison with LLM self-explanations, and Yannis Katsis for his advice on the setup for this experiment. The work of Lucas Monteiro Paes was supported by the Apple Scholars in AI/ML Fellowship.
Appendix A More on Related Work
Table 3 compares the features of the proposed MExGen framework to alternative perturbation-based explanation methods. The remainder of this appendix discusses other categories of explanation methods.
Hierarchical Explanations
A line of work Singh et al. (2019); Jin et al. (2020); Chen et al. (2020); Ju et al. (2023) has developed hierarchical explanations for sequence models including LLMs, which can reveal compositional interactions between words and phrases. In particular, the HEDGE algorithm of Chen et al. (2020) was identified by Mosca et al. (2022) as “arguably the most suitable choice” for NLP input attribution, in part because it builds its hierarchy in a top-down, divisive fashion (as opposed to bottom-up agglomeration Singh et al. (2019); Ju et al. (2023)), which is more practical for long texts. However, HEDGE is specific to classification because it measures feature importance based on classification margin.
Gradient-Based Explanations
Gradient-based methods provide input attribution explanations Simonyan et al. (2013); Sundararajan et al. (2017); Shrikumar et al. (2017); Bach et al. (2015), but they require access to model gradients with respect to the input.
Attribution for Generative Language Models.
Concurrently with our work, other studies have explored attribution techniques for generative LMs Li et al. (2024); Cohen-Wang et al. (2024); Sarti et al. (2023). GiLOT Li et al. (2024) employs optimal transport to compute explanations. However, it focuses on the overall distributional changes across all possible outputs rather than attributing specific generated outputs to particular inputs, as MExGen does.
ContextCite Cohen-Wang et al. (2024) extends LIME Ribeiro et al. (2016) to generative models by considering parts of the context as features. While effective, ContextCite operates at a single predefined granularity level (e.g., sentence level), lacking the multi-level nature of MExGen, which improves explanation performance.
Inseq Sarti et al. (2023) is a toolkit for attribution of generative LMs, offering methods and visualization tools for context attributions. Its methods focus on computing the importance of each input token for every token in the generated output. However, this approach may not fully capture the influence of larger units in the input, and it also does not explain the entire output.
A key distinction of MExGen is its ability to function without access to the model’s logits by utilizing text-only scalarizers, whereas GiLOT, ContextCite, and Inseq all assume access to these logits.
Self-Explanations & Numerical Attributions
Other works have used the generative LM itself to provide explanations in line with subsequent outputs.
An example of self-explanation is the chain-of-thought (CoT) Wei et al. (2022), where the model explicitly generates the reasoning in natural language used to produce its output. However, CoT has issues with stability Turpin et al. (2023) and faithfulness Atanasova et al. (2023); Lanham et al. (2023) in addition to significant variation in quality Nagireddy et al. (2024). Moreover, evaluating the faithfulness of CoT is challenging, and Parcalabescu and Frank (2024) argued that existing faithfulness evaluations actually measure self-consistency. Parcalabescu and Frank (2024) proposed their own measure of self-consistency called CC-SHAP, which adapts SHAP in a way similar to how we adapt P-SHAP. CoT is less relevant to our work since it provides explanations in natural language rather than numerical attributions.
Another approach to self-explanation is prompting the model to explain its predictions (e.g., by ranking the most important parts of its context) Camburu et al. (2018); Huang et al. (2023b); Kroeger et al. (2024); Madsen et al. (2024). However, prior work has found that these self-explanations may be less faithful when used for in-context classification tasks Madsen et al. (2024); Huang et al. (2023b).
Recently, Fragkathoulas and Chlapanis (2024) employed a modified version of MExGen to identify words in the input context that were essential for the model’s prediction in a question-answering task. Then, Fragkathoulas and Chlapanis (2024) compared the keywords they identified with the keywords extracted using self-explanation, showing that self-explanations are not as faithful as their method. This is a first step in demonstrating that self-explanations are not as faithful as numerical attributions, as provided by MExGen. Our work generalizes theirs to additional generative tasks and to self-explanations in the form of rankings.
Rationalization
Yet another line of self-explanation methods falls under the class of rationalization methods. Lei et al. (2016) propose to simultaneously train a rationale extractor with a predictor. The rationale extractor essentially selects parts of the text to use for prediction that fulfill two criteria: they are interpretable and maintain nearly the same prediction as using the full text. As these parts of the text are interpretable, they offer an explanation for the final prediction. Lei et al. (2016) employed regularizations to keep the rationales short while encouraging contiguous text, and a sequence of recent literature has focused on improving these rationales.
Liu et al. (2022) propose to share the encoder parameters of the extractor and predictor, with the intuition that the extractor and predictor are both seeking to find the most informative part of the input text. Liu et al. (2023b) takes a different approach and considers removing what is unimportant rather than selecting what is important; they minimize a criterion for dependence of the prediction on the text which essentially detects what parts of the text is not required for prediction. Liu et al. (2023c) use a theoretical analysis of how the Lipschitz constant of the predictor affects the extractor to propose how to control the learning rate for obtaining better rationales. Liu et al. (2023a) introduce an ensemble of extractors (multiple extractors with different initializations) into the rationalization training framework and find improved prediction performance. At inference time, only the first extractor of the ensemble is used for interpretability. Liu et al. (2024b) further include a predictor on the full text into a regularization term in order to better align predictions on extracted text with those of the full text. Most recently, Liu et al. (2024a) build on Liu et al. (2023b) and propose a loss criterion that attempts to treat spurious correlations in the text as if it were noise.
Appendix B Further Details on MExGen
This appendix provides more details on the MExGen framework that are of a more general nature. For parameter settings and other details specific to our experiments, please refer to Appendix C.
B.1 Scalarizers
@@
The inputs $x$ in our case are perturbed versions of the original input $x^{o}$. Some of these perturbations however may be significantly different semantically and cause the probability of generating the target output $y^{o}$ to decrease dramatically. For this reason, the Log Prob scalarizer may require access to logits for improbable tokens conditioned on $x$, which can be ensured if logits are available for the entire vocabulary, but not if only the top $k$ logits are provided.
@@
BARTScore
BARTScore is the same as (1) except with $y$ in place of $x$ and an auxiliary LM $f_{\textit{BART}}$ (not necessarily a BART model) in place of the LM $f$ being explained. It thus measures the probability of $f_{\textit{BART}}$ generating $y^{o}$ given $y$ as input.
@@
We included this scalarizer to represent our understanding of how the abstractive summarization example SHAP (2024b) in the SHAP library handles the text-only API case. We investigated the code behind this SHAP example, and specifically the TeacherForcing class that it uses. Our understanding of the code is that it obtains proxy log-odds for tokens in the target output $y^{o}$ by taking log-odds from an auxiliary summarization model $f_{\textit{SUMM}}$, with input $y$ and output set to $y^{o}$. If we then average the log-odds over the tokens in $y^{o}$ (similar to (1)), the result is similar to the BART scalarizer with $f_{\textit{BART}}=f_{\textit{SUMM}}$ (with a possible discrepancy between log-odds versus log probabilities). Our experiments show that the BART and SUMM scalarizers are indeed very similar.
@@
This scalarizer is based on the intuition that Log NLI entailment is a kind of similarity. We use the NLI model to predict the log-odds of entailment given $y^{o}$ as premise and $y$ as hypothesis, and optionally in the other direction as well. If both directions are used, we take the geometric mean of the two entailment probabilities and then convert back to log-odds. We note that bi-directional entailment has been used to approximate semantic equivalence, for example in Kuhn et al. (2023), but here we use the log-odds scores and not just the predicted labels.
B.2 Phrase Segmentation
Here we describe the phrase segmentation algorithm mentioned in Section 3.2. The algorithm starts with the root token of the tree and checks whether each child subtree of the root is shorter (in terms of tokens) than a maximum phrase length parameter. If it is shorter, then the subtree constitutes a phrase, and if it is not, the algorithm is recursively applied to the subtree. The root token of each (sub)tree is also taken to be a phrase. Once the sentence has been recursively segmented into phrases in this manner, a second pass is performed to re-merge some phrases that have become too fragmented, thus controlling the number of phrases which is desirable for computation and interpretation. Specifically, phrases that constitute noun chunks (as identified by spaCy) are merged, and certain single-token phrases are merged with their neighbors. Further notes:
@@ B.2 Phrase Segmentation
B.3 L-SHAP
This local variation of SHAP was proposed by Chen et al. (2019) to decrease the number of model inferences relative to SHAP, which requires exponentially many inferences. In our context, L-SHAP does so by only perturbing units that are within a constant-size neighborhood of the current unit being attributed to. This makes the number of inferences scale linearly with the number of units. More precisely, for unit of interest $s\in[d]$, we consider only the radius-$M$ neighborhood $\mathcal{N}_{s}^{M}=\{s-M,\dots,s-1,s+1,\dots,s+M\}$ (truncated to $1,\dots,d$ if necessary) and subsets of $\mathcal{N}_{s}^{M}$ with cardinality up to $K$. Then the attribution score $\xi_{s}$ for unit $s$ is given by
$\xi_{s}=\frac{1}{K+1}\sum_{A\subseteq\mathcal{N}^{M}_{s}:|A|\leq K}\binom{|\mathcal{N}^{M}_{s}|}{|A|}^{-1}\\ \times\left(S(x^{(A)};y^{o},f)-S(x^{(A\cup\{s\})};y^{o},f)\right),$ (3)
where $x^{(A)}$ is a perturbation of $x^{o}$ in which units $j\in A$ are perturbed (just dropped in our work).
Appendix C Automated Evaluation Details
This appendix documents choices and parameter settings used in the experiments in Section 4. We follow the order of presentation in Section 4.1: datasets and LM inference, choices for the MExGen framework (Appendix C.3), baseline attribution methods (Appendix C.4), perturbation curve details, and computing environment.
C.1 Datasets
The GitHub repository [^6] https://github.com/EdinburghNLP/XSum that enables the XSUM dataset Narayan et al. (2018) to be rebuilt is made available under the MIT license. The CNN/DM dataset See et al. (2017) is made available by HuggingFace [^7] https://huggingface.co/datasets/abisee/cnn_dailymail under the Apache-2.0 license. Both datasets are intended for abstractive summarization, which is how we use them. Both datasets consist of news articles and thus contain the names of individuals. Since these names were reported by the media in the original articles, they are already in the public domain and there is no further issue.
SQuAD is distributed under the CC BY-SA 4.0 license. The dataset is intended for evaluating reading comprehension by answering questions on the contexts in the dataset, which is how we use it as well. Since the contexts are taken from Wikipedia articles, some of them name individuals, but again these articles are already in the public domain. We selected 1000 validation set examples at random from SQuAD because selecting the first 1000 examples yielded insufficient diversity.
C.2 Language model inference
DistilBART and Flan-T5-Large
The two models that we treated as providing full logit access, DistilBART and Flan-T5-Large, were downloaded from HuggingFace under the Apache-2.0 license. The DistilBART model was trained on the summarization datasets XSUM and CNN/DM, and we use it for summarization. Flan-T5-Large is a general-purpose LM intended for research on LMs, consistent with our use.
The models were called through their .generate method. For generating the original (target) output (corresponding to the original input), max_new_tokens was set to None (i.e., the default). When using the text-only scalarizers, perturbed output texts (corresponding to perturbed inputs) are also generated, and for these, max_new_tokens was set to $1.5$ times the number of tokens in the target output. The Log Prob scalarizer computes the log probability of generating the target output, so max_new_tokens is not needed in this case. All other hyperparameters were left at default settings (for example greedy decoding was used).
Flan-UL2 and Llama-3
Flan-UL2 and Llama-3-8B-Instruct are general-purpose LMs. Flan-UL2 is distributed by HuggingFace under the Apache-2.0 license while Llama-3 is distributed under its own Meta Llama 3 Community License. [^8] https://www.llama.com/llama3/license/ We however accessed Flan-UL2 and Llama-3 using an LLM API service provided by IBM (IBM, 2024) (no longer in existence).
For these API calls to Flan-UL2 and Llama-3-8B-Instruct, we used greedy decoding and max_new_tokens $=100$. In the case of Llama-3, we used its chat template and provided the following system prompts for summarization and QA respectively: (summarization) “Summarize the following article in one sentence. Do not preface the summary with anything.” (QA) “Please answer the question based on the provided context. Answer with a short phrase or sentence.” No chat template or system prompt was necessary to prompt Flan-UL2 to summarize.
DeepSeek-V3 and Granite-3.3
The DeepSeek-V3 and Granite-3.3-8B-Instruct models used in the self-explanation experiment of Section 4.4 were called using a different LLM API service provided by IBM. Similar to Llama-3, we used greedy decoding, max_tokens $=100$, the model’s chat template, and the same system prompt for summarization above. We also fixed the LLM’s random seed so that it would generate the same output given the same input. For generating self-explanations in the form of a list of tags, max_tokens was increased to 500 to accommodate long lists. For Granite-3.3, we were able to obtain log probabilities by setting max_tokens $=0$ and logprobs $=0$, which returns log probabilities of tokens in the input given to the LLM.
C.3 MExGen
@@
  • Sim”: We use the all-MiniLM-L6-v2 embedding model from the SentenceTransformers package Reimers and Gurevych (2019). [^9] This model can be found at https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 and is distributed under the Apache-2.0 license.
  • BERT”: We use the model deberta-v2-xxlarge-mnli [^10] This model can be found at https://huggingface.co/microsoft/deberta-v2-xxlarge-mnli under the MIT license. He et al. (2021) to compute BERTScore Zhang et al. (2020) (MIT license). This is the same model that we use for the Log NLI scalarizer. Our initial reason for doing so was to see whether the two scalarizers would be very similar because of this choice (they are not as they operate on different principles). We take the “F1-score” output as the BERTScore.
  • BART” and “SUMM”: For “SUMM”, we follow SHAP (2024b) in using the same distilbart-xsum-12-6 model as both the scalarizing summarization model as well as the primary summarization model to explain. For “BART”, i.e., BARTScore Yuan et al. (2021), we use the code [^11] https://github.com/neulab/BARTScore, Apache-2.0 license from the authors and also instantiate it with the distilbart-xsum-12-6 model. The purpose was to determine whether the “BART” and “SUMM” scalarizers are very similar when instantiated with the same model, which is indeed the case.
@@
In the case where log probabilities are not available from the LLM being explained, the above text-only scalarizers contribute to the overall computational cost of MExGen, but only in a secondary way. This is because our choices of scalarizers are small LMs (at most $1.5$B parameters for deberta-v2-xxlarge-mnli). The Sim scalarizer was the most computationally efficient because it uses a SentenceTransformer model. The other scalarizers were more comparable to each other, with the Log NLI scalarizer being the most costly because of our choice to compute bi-directional entailment (two inferences versus one).
@@
As mentioned in Section 3.2, we use spaCy v3.6 Honnibal et al. (2020) (distributed under the MIT license) for sentence and word segmentation. Our custom phrase segmentation algorithm has one main parameter, the maximum phrase length, which we set to $10$ spaCy tokens (not counting spaces and punctuation).
In terms of units that are not of interest for attribution, we exclude non-alphanumeric units (generally punctuation and newlines) and prompt template elements, for example start-of-message and end-of-message tokens like “<|start_header_id|>user<|end_header_id|>”, prefixes such as “Context: ”, and system prompts. For QA, we exclude the question and allow only the context to be attributed to.
@@
The decision here is how many top sentences are refined into phrases in the summarization experiments and into words in the QA experiments. For this we follow Algorithm 1 with the following parameter settings: For summarization with the DistilBART model, the number of top sentences is set to $k=3$ and the significance threshold is set to $\phi=1/3$ for sentence-level scores from C-LIME and LOO and $\phi=0.3$ for L-SHAP. For summarization with Flan-UL2, Llama-3, DeepSeek-V3, and Granite-3.3, we set $k$ to correspond to the top 25% of sentences, rounding to the nearest integer and ensuring $k\geq 1$. A significance threshold is not used, i.e., $\phi=-1$. For QA, we simply take the top $k=1$ sentence as the context paragraphs in SQuAD tend to have only a handful of sentences, and again do not use a significance threshold ($\phi=-1$).
You provided the summary below of an article, also below. The article is divided into units (sentences or phrases), which are numbered in the format: <u0> unit 0 <u1> unit 1 … Please list the {top_k} units that were most important for you to produce this summary. List them in order from most important to least important. List only the unit numbers, for example “<u3>, <u1>, <u4>”. Summary: {summary} Article: <u0> {units[0]} <u1> {units[1]} …
Figure 6: Prompt for LLM self-explanation
C-LIME
For C-LIME, the parameters controlling the perturbations are the constant of proportionality between the number of perturbations $n$ and number of units $d$, and the maximum number of units $K$ perturbed at one time. For the smaller DistilBART and Flan-T5-Large models, $n/d$ is set to $10$ and $K$ to $3$, except for sentence-level attributions on SQuAD where $K=2$.
For Flan-UL2, Llama-3-8B-Instruct, DeepSeek-V3, and Granite-3.3-8B-Instruct, since inference is more costly for these larger models, we take the following approach. First, at the sentence level of attribution, we only use LOO to identify the top sentences and do not use C-LIME. At the phrase level for the summarization datasets, we use C-LIME with $K=2$ and $n$ equal to the total number of phrases (in the entire document, not just in the sentences selected for refinement). Similarly, at the word level for SQuAD, we set $K=2$ and $n$ equal to the total number of words. In both cases this corresponds to a ratio $n/d\approx 5$.
C-LIME has two additional aspects to consider:
Sample weighting: ($\pi(z)$ in (2)) Since we limit the number of units $K$ that are simultaneously perturbed to a small integer, we also no longer use LIME’s sample weighting scheme. Instead, we give each subset cardinality $k=0,\dots,K$ the same total weight (say $1$ without loss of generality), and then distribute this weight uniformly over the subsets of that cardinality.
Regularization: Different regularizers $R(w)$ can be used in (2), e.g. $\ell_{2}$ or $\ell_{1}$. In our experiments however, we do not regularize (i.e., $R(w)\equiv 0$) and compute a fully dense solution. This allows all units to be ranked to facilitate evaluation of perturbation curves.
@@
For L-SHAP, the two main parameters are the local neighborhood radius $M$ and the maximum number of neighbors $K$ perturbed at one time. (Note that $K$ does not include the unit of interest, so altogether the maximum number of perturbed units is $K+1$.) For the DistilBART and Flan-T5-Large models, we take $M=2$ and $K=2$. For Flan-UL2, Llama-3, DeepSeek-V3, and Granite-3.3, we again use only LOO at the sentence level and do not use L-SHAP. For phrase-level or word-level L-SHAP, we set $M=1$ and $K=2$.
C.4 Baseline attribution methods